# Quantum property testing in sparse directed graphs

### v1

@@ 2.3 Property testing
In decision problems, the algorithm has to decide if the input satisfies a property or not. In the case of property testing the question is relaxed: the algorithm has to distinguish inputs that satisfy the property from those that are “far” (according to some distance measure) from any input that satisfies it.
@@
Applying this to the case of bounded-degree graphs degree bound $d$ and query access to the adjacency list, the distance of two graphs is the number of edges where they differ divided by $|V|d$. The distance of an object $x$ from a property $P$ is the minimum distance between $x$ and any object that satisfies $P$.
@@ 2.4 Problem definitions
We now define the problems we study and argue about certain relations between them. While the problems are phrased as decision problems, ultimately we will care about the quantum query complexity for testing the corresponding properties. The complexity is going to be parameterised by a parameter $k$. Moreover, the parameter $k$, the degree bound $d$ and the parameter $\varepsilon$, are all considered to be constants throughout this paper.
@@
Let $G=(V,E)$ be a digraph. A set $S\subseteq V$ is called a source component if it induces a strongly connected subgraph in $G$, and in $G$ there is no edge from $V\setminus S$ to $S$.
@@
We will now state the decision variant of several problems. The “property” corresponding to a decision problem is the set of inputs that should be accepted in the decision problem.
@@
For proving a lower bound, we will look at a special case of the main problem: $k$-star-freeness. Notice that a $k$-star has $k$ source components, hence a lower bound for this problem implies the same lower bound for the more general $k$-source-subgraph-freeness problem.
@@
For the lower bound on $k$-star-freeness testing, we are going to use as a “helper problem” the decision variant of the $k$-collision problem.
@@
Accept iff $s$ is $k$-collision-free, i.e. there is no $i_{1},\dots,i_{k}\in[N]$ with $s_{i_{1}}=\dots=s_{i_{k}}$
@@ 3 Quantum algorithm for testing subgraph-freeness
In this section, we prove that there is a quantum speedup for testing $H$-freeness in directed graphs with $d$-bounded out-degree, for any graph $H$ that has $k$ source components. For large but constant $k$, the speedup is near-quadratic. The algorithm relies on the following simple claim. It shows that if $G$ is far from being $H$-free then it contains many (not necessarily disjoint) copies of $H$. This problem was studied in [goldreich2002property] in the classical setting.
@@
Assume that an $N$-vertex graph $G$ with $d$-bounded out-degree is $\varepsilon$-far from $H$-freeness. Then for each source component of $H$, there are $\Omega(\varepsilon N)$ vertices that are part of that component in an $H$-subgraph in $G$.
@@
Since $G$ is $\varepsilon$-far from $H$-freeness, more than $\varepsilon Nd$ edges need to be removed to get an $H$-free graph. Thus, for each of the $k$ source components of $H$, there are more than $\varepsilon Nd$ edges in $G$ for which there is an $H$-subgraph where the edge starts from a vertex of the source component. Otherwise, deleting these (at most $\varepsilon Nd$) edges would make that source component disconnected from the rest of $H$ in all the $H$-instances, making $G$ free from having $H$ as a subgraph. This, in turn, implies that for each of the source components of $H$, there are at least $\varepsilon N$ vertices that are part of the source component in some $H$-subgraph in $G$, since every vertex can have at most $d$ outgoing edges. ∎
To illustrate the algorithm, we first consider the $k=2$ case to build our intuition. Our algorithm takes inspiration from the BHT algorithm for collision finding [BHT]. In the following, we set $m=|V(H)|$, number of vertices in $H$. We use the shorthand BFS for breadth-first search.
@@
    Randomly sample a vertex subset $\mathcal{S}$ in $G$ of size $t=\Theta(N^{1/3})$. Perform a depth-$m$ BFS from every vertex $v\in\mathcal{S}$.
@@
    Perform a Grover search over the remaining vertices $V\backslash\mathcal{S}$ to look for a vertex that – after doing a depth-$m$ BFS from it – completes an $H$-instance with one of the sampled vertices.
@@
Note that if $G$ is $H$-free, then the above algorithm will always accept. Now we need to argue that if $G$ is $\varepsilon$-far from being $H$-free, then with constant probability the above algorithm will find a copy of $H$ and thus reject.
To do so, pick one source component of $H$. By Proposition 3.1, with high probability a constant fraction of the $t$ vertices in $\mathcal{S}$ are part of that source component in an $H$-subgraph in $G$. For such vertices, the BFS in step 1 will discover the entire source component, as well as all other vertices reachable from that source component in $H$. Then, in step 2 we effectively search for a vertex that is in the remaining source component of such an instance of $H$ that we already partly discovered. Again by Proposition 3.1, with high probability there are $\Omega(t)$ many of them. By doing a BFS from that vertex, we uncover the resulting $H$-subgraph. This proves correctness of the algorithm.
Finally, we bound the algorithm’s query complexity. Step 1 only makes $O(t)=O(N^{1/3})$ many queries. For step 2, we argued that there are $\Omega(t)$ many “good” vertices, and hence Grover search will make $O(\sqrt{N/t})=O(N^{1/3})$ quantum queries.
@@
In order to extend the $k=2$ case described above to larger $k$, we first try to find many partial $H$-instances with $k-1$ source components found, and then extend one of them to a complete $H$-instance. We present a brief description of our algorithm below, where $m$ is the number of vertices of $H$:
@@
    Randomly sample a vertex subset $\mathcal{S}_{1}$ in $G$ of size $t_{1}$. Perform a depth-$m$ BFS from every vertex $s\in\mathcal{S}_{1}$.
@@
      Perform a Grover search $t_{i}$ times on the vertices $V\setminus\mathcal{S}_{i-1}$ to find a set $\mathcal{S}_{i}$ of $t_{i}$ many partial $H$-instances that cover $i$ of its source components.
@@
      Set $\mathcal{S}_{i+1}=\mathcal{S}_{i-1}\cup\mathcal{S}_{i}$.
@@
    Perform Grover search on $V\backslash\mathcal{S}_{k}$ to find a complete $H$-instance.
@@
The correctness proof is same as the $k=2$ case, and is omitted. To bound the query complexity, note that the first Grover search finds $t_{2}$ partial $H$-instances with $2$ of its source components found, which takes $O(t_{2}\sqrt{N/t_{1}})$ queries. Similarly, for $i$-th iteration, the algorithm performs $O(t_{i}\sqrt{N/t_{i-1}})$ quantum queries for every $i\in[k-1]$. Finally, finding one complete $H$-instance costs $O(\sqrt{N/t_{k-1}})$ queries. Thus the total query complexity is $O(t_{1}+\sum_{i=1}^{k-1}t_{i+1}\sqrt{N/t_{i}})$ with $t_{k}=1$. Similar to the multi-collision algorithm in [zhandry_collision, Section 3], we can equate all terms by setting $t_{i}=\Theta\left(N^{\frac{2^{k-i}-1}{2^{k}-1}}\right)$, which yields the claimed quantum query complexity $O\left(N^{\frac{1}{2}\left(1-\frac{1}{2^{k}-1}\right)}\right)$. ∎
4 Quantum lower bound
@@
Let $k\geq 3$ and $0<\varepsilon<1/(4^{k-1}\lceil 20(2k)^{k/2}\rceil)$ be constants. Let $N$ be a large enough positive integer. Then the quantum query complexity of the $\varepsilon$-testing of $k$-collision-freeness of a sequence of integers $S=(s_{1},\dots,s_{N})\in[N]^{N}$ with parameter $\varepsilon$ is $\Omega(N^{1/2-1/(2k)}/\ln^{2}N)$.
@@ 4.1 The (dual) polynomial method
We consider a property on Boolean vectors as a function $f:D\subseteq\{-1,1\}^{n}\to\{-1,1\}$. Since the work of [beals_poly] it has been known that the acceptance probability $p(x)$ of a $T$-query bounded-error quantum algorithm on input $x\in D$ is a polynomial of degree at most $2T$. Thus, the polynomial $1-2p(x)$ must be a good approximation of $f$. Observe that $1-2p(x)$ remains bounded outside $D$ since $p(x)$ remains a probability defined by the algorithm, with no constraint.
@@
Let $f:D\subseteq\{-1,1\}^{n}\to\{-1,1\}$ and $\delta>0$. If a quantum algorithm computes $f$ on $D$ with error $\delta$ using $T$ queries then there is a polynomial $p$ of degree at most $2T$ that $2\delta$-approximates $f$ on $D$.
@@
We now consider a property of a sequence of integers as a function $F:D\subseteq[R]_{0}^{N}\to\{-1,1\}$. The symbol 00 will play a special role that will be exhibited later on. Unfortunately one cannot just take the polynomial of those integers. The standard approach (see [AAronson02]) is to encode $s=(s_{1},\dots,s_{N})\in[R]_{0}^{N}$ into binary variables $x=(x_{i,j})_{i\in[N],j\in[R]_{0}}\in\{-1,1\}^{N(R+1)}$ encoding whether $s_{i}=j$ as follows: $x_{i,j}=-1$ if $s_{i}=j$, and $x_{i,j}=1$ otherwise. Let $H_{b}^{N(R+1)}\subseteq\{-1,1\}^{N(R+1)}$ be the set of all possible encodings of vectors $s$, that is for every $i\in[N]$ there is exactly one $j\in[R]_{0}$ such that $x_{i,j}=-1$.
@@
But before doing this, we are going to relax the constraints on the domain $D_{b}$ in the case of symmetric functions, while we decrease its dimension. When $F$ is symmetric (i.e. $F(s)=F(s\circ\pi_{N})$ for any permutation $\pi_{N}$ of $[N]$), one can instead define a function $F_{\leq N}$ with weaker constraints by removing the variables corresponding to the symbol 00. Define $H_{\leq N}^{NR}$ as the set of length-$(NR)$ binary vectors with Hamming weight at most $N$. Given any $x\in H_{\leq N}^{NR}$, we define its frequency vector $z(x)=(z_{0},z_{1},\dots,z_{R})$ with $z_{j}=\#\{i:x_{ij}=-1\}$, for $1\leq j\leq R$, and $z_{0}=N-z_{1}-\ldots-z_{R}$. From the vector $z(x)$, one can define a valid sequence of integers $s(x)\in[R]_{0}^{N}$: it can be any sequence from $[R]_{0}^{N}$ that has frequency vector $z(x)$. Now we can define $F_{\leq N}$ on domain $D_{\leq N}$ as
@@
Lemma 4.7 ([AAronson02, ambainis_poly],[polynomial_strikes_back, Theorem 6.5]).
Let $F:D\subseteq[R]_{0}^{N}\to\{-1,1\}$ be symmetric and $\delta>0$. Define $H_{\leq N}^{NR}\subseteq\{-1,1\}^{NR}$ and $F_{\leq N}:D_{\leq N}\subseteq H_{\leq N}^{NR}\to\{-1,1\}$ as above. If a quantum algorithm computes $F$ on $D$ with error $\delta$ using $T$ queries then there is a polynomial $p$ of degree at most $2T$ that double-promise $2\delta$-approximates $F_{\leq N}$ on $D_{\leq N}$.

### v2

@@ 2.3 Property testing
In total decision problems, the algorithm has to decide if the input satisfies a property or not. In the case of property testing, the question is relaxed: the algorithm has to distinguish inputs that satisfy the property from those that are “far” (according to some distance measure) from any input that satisfies it.
@@
Applying this to the case of bounded-degree graphs with degree bound $d$ and query access to the adjacency list, the distance of two graphs is the number of edges where they differ divided by $|V|d$. The distance of an object $x$ from a property $\mathcal{P}$ is the minimum distance between $x$ and any object that satisfies $\mathcal{P}$.
@@
2.4 Grover search
We are not going into details about quantum computing, because for this paper, it suffices to state very few results in the field and use them in a black box way. One of the most important results in quantum computing is Grover’s algorithm for unordered search [Grover]: finding a marked element in an unordered database of size $N$ takes $\Theta(N)$ queries classically, but the quantum query complexity of the same task is $\Theta(\sqrt{N})$. We are going to use a particular variant of this result that has been used many times in the literature (see e.g., [Ambainis_search, Item 3 in Section 2.2], which was implicitly proved in [BBHT_search]).
Theorem 2.2.
Let $1\leq t_{0}\leq N$. There exists a quantum algorithm that, given $t_{0}$ and query access to any function $f:S\to\{0,1\}$, makes $O(\sqrt{N/t_{0}})$ queries to $f$ and outputs either “not found” or an element uniformly at random in $f^{-1}(1)$. Moreover, when $|f^{-1}(1)|\geq t_{0}$, the latter occurs with high constant probability.
Remark 2.3.
In practice, we will use this theorem when querying $f(z)$, for $z\in S$, requires making $c$ queries to an input graph $G$ with vertex set $S$. In that case the total query complexity to $G$ is $O(c\sqrt{N/t_{0}})$.
@@ 2.5 Problem definitions
We now formally define the problems we study and argue about certain relations between them. While the problems are phrased as decision problems, ultimately we will care about the quantum query complexity for testing the corresponding properties. The complexity is going to be parameterized by a parameter $k$. Moreover, the parameters $k,\varepsilon$ and the degree bound $d$ are all considered to be constants throughout this paper.
@@
Let $H=(V,E)$ be a digraph. A set $S\subseteq V$ is called a source component if it induces a strongly connected subgraph in $H$, and there is no edge from $V\setminus S$ to $S$ in $H$.
@@
Notice that a $k$-star has $k$ source components each consisting of a single vertex.
We will now state the decision variants of several problems. The “property” corresponding to a decision problem is the set of inputs that should be accepted in the decision problem.
@@
For proving a lower bound, we will look at a special case of the main problem: $k$-star-freeness. Notice that since a $k$-star has $k$ source components, a lower bound for this problem implies the same lower bound for the more general $k$-source-subgraph-freeness problem.
@@
For the lower bound on $k$-star-freeness testing, we are going to use as a “helper problem” the testing variant of the $k$-collision problem.
@@
Accept iff $s$ is $k$-collision-free, i.e., there is no $i_{1},\dots,i_{k}\in[N]$ with $s_{i_{1}}=\dots=s_{i_{k}}$
@@ 3 Quantum algorithm for testing subgraph-freeness
In this section, we prove that there is a quantum speedup for testing $H$-freeness in directed graphs with $d$-bounded out-degree, for any graph $H$ that has $k$ source components. For large but constant $k$, the speedup is near-quadratic. This problem was studied in [goldreich2002property] in the classical setting. Our algorithm can be seen as a generalization of the one in [zhandry_collision] to graphs. Let us start with the definition of source-disjointness which will be used in the analysis of our algorithm.
Definition 3.1 (Source-disjointness).
Let $G$ be a directed graph such that it contains two subgraphs $H_{1}$ and $H_{2}$. We say that $H_{1}$ and $H_{2}$ are source-disjoint if the union of the source components of $H_{1}$ is disjoint from the union of the source components of $H_{2}$.
Moreover, we need to prove the following simple proposition. It shows that if $G$ is far from being $H$-free, then it contains many source-disjoint copies of $H$, that is, copies of $H$ that are source-disjoint subgraphs of $G$.
@@
Let $H$ be an $h$-vertex graph with $k$ source components. Assume that a $d$-bounded out-degree directed graph $G$ on $N$ vertices is $\varepsilon$-far from $H$-free. Then $G$ contains at least $\varepsilon N/h=\Omega(N)$ source-disjoint copies of $H$.
@@
We prove the result by contraposition. Consider a maximal set $M$ of source-disjoint copies of $H$ in $G$, and assume that $|M|\leq\varepsilon N/h$. Let $U$ denote the union of all the vertices in the source components of the copies in $M$. This implies that if one deletes all the outgoing edges of all the vertices in $U$, then $G$ becomes $H$-free. Indeed, if there remained an $H$-copy then all its source components are disjoint from $M^{\prime}$ (as in a source component every vertex has at least one outgoing edge), contradicting the fact that $M$ was maximal.
Since $|U|\leq|M|\cdot h$, the number of those deleted edges is at most $|U|\cdot d\leq|M|\cdot hd\leq\varepsilon Nd$. Therefore, the resulting graph is both $H$-free and $\varepsilon$-close to the original graph $G$. This proves the contraposition of the proposition. ∎
We will use Breadth-first search (BFS) in order to explore $G$ layer by layer: starting from a given vertex, first it explores its direct (out-)neighbors, then their unexplored (out-)neighbors etc. We will run BFS up to some limited depth $\ell$, so that the $-\ell$ BFS algorithm has query complexity at most $d^{\ell}$ where $d$ is the maximum (out-)degree of $G$. In our application, $d$ and $\ell$ are constants, and thus the query complexity is $d^{\ell}=O(1)$.
3.1 The algorithm for $k=2$
To illustrate our algorithm, we first consider the $k=2$ case to build some intuition. Here, our algorithm generalizes the BHT algorithm for collision finding [BHT] in the context of graphs. The high level idea is that if we manage to sample a vertex from each of the two source components of an $H$-subgraph (a collision), then by querying their “surroundings” we will discover the $H$-instance. In the following, we set $h=|V(H)|$, the number of vertices in $H$.
@@ 3.1 The algorithm for $k=2$
[…]
@@
    Perform Grover search over the remaining vertices $V\backslash\mathcal{S}$ in the following way. A vertex $v$ is marked if there exists another vertex $u\in\mathcal{S}$ such that $u$ and $v$ are from the 2 different source components of an $H$ subgraph of $G$.
@@
Note that if $G$ is $H$-free, then the above algorithm will always accept. Now we need to argue that if $G$ is $\varepsilon$-far from being $H$-free, then with constant probability, the above algorithm will find a copy of $H$ and thus will output reject.
[…]
3.2 The algorithm for general $k$
@@
In order to extend the $k=2$ case described above to larger $k$, we first try to find many partial $H$-instances with $k-1$ source components found, and then extend one of them to a complete $H$-instance. We present a brief description of our algorithm below, where $h$ is the number of vertices of $H$:
@@
[…]
@@
      Perform a Grover search $t_{i}$ times on the vertices $V\setminus\mathcal{S}_{i-1}^{\prime}$ in the following way. A vertex $v$ is marked if there exist $i-1$ other vertices $u_{j}\in\mathcal{S}_{j}$ for each $j\in[i-1]$ such that $u_{1},\dots,u_{i-1}$ and $v$ are from $i$ different source components of an $H$ subgraph of $G$. If we do not find $t_{i}$ vertices like this, output Reject, otherwise let $\mathcal{S}_{i}$ denote the set of the vertices $v$ that we found.
@@
      Set $\mathcal{S}_{i}^{\prime}=\mathcal{S}_{i-1}^{\prime}\cup\mathcal{S}_{i}$.
@@
    Perform Grover search on $V\backslash\mathcal{S}_{k-1}^{\prime}$ to find a complete $H$-instance. I.e., a vertex $v$ is marked if there exist $k-1$ other vertices $u_{j}\in\mathcal{S}_{j}$ for each $j\in[k-1]$ such that $u_{1},\dots,u_{k-1}$ and $v$ are from the $k$ different source components of an $H$ subgraph of $G$.
@@
The correctness proof is similar to the $k=2$ case. Proposition 3.2 tells us that in $\mathcal{S}_{1}$ there are $\Omega(t_{1})$ many vertices that are from a source component of an $H$-copy. Because of the source-disjointness of the $H$-copies, when $i=2$, there are $\Omega(t_{1})$ many 1-partial solutions that can be extended to a complete $H$ instance by disjoint remaining source components. As Grover search provides uniformly random marked elements, a constant fraction of the $t_{2}$ many 2-partial solutions are actually extendable to $H$ in a similar, disjoint way. This continues to be true in each iteration: (with high probability) a constant fraction of the $t_{i-1}$ many $(i-1)$-partial solutions are extendable to complete $H$ instances by disjoint remaining source components. This way, the last step is going to find an $H$-subgraph with high probability.
[…]
We note that there is no need for a $\textnormal{polylog}(N)$ factor in the query complexity, which could come from a commonly used way to boost up the success probability of Grover’s algorithm. This stems from two observations. First, consider the case where $K$ among $N$ elements are marked, with a given lower bound $L\leq K$, and we wish to find $R\leq L$ such elements. If $R\ll L$, then (say) $100R$ repetitions of Grover should return at least $R$ marked elements with probability at least $2/3$ while making $O(R\sqrt{N/L})$ queries, without extra log-factors. This is due to the fact that one can simply ignore any unsuccessful Grover runs. In our case we set $R=t_{i+1}\ll L=t_{i}$. Finally, since there are $k$ iterations in the algorithm and $k$ is constant, a factor of $\log k$ would not add up to the query complexity of our algorithm in terms of $N$. ∎
4 Collision-freeness lower bound
@@
Let $k\geq 3$ and $0<\varepsilon<1/(4^{k-1}\lceil 20(2k)^{k/2}\rceil)$ be constants and $N$ be a large enough positive integer. Then the quantum query complexity of the $\varepsilon$-testing of $k$-collision-freeness of a sequence of integers $S=(s_{1},\dots,s_{N})\in[N]^{N}$ with parameter $\varepsilon$ is $\Omega(N^{1/2-1/(2k)}/\ln^{2}N)$.
@@ 4.1 The (dual) polynomial method
We consider a property on Boolean vectors as a function $f:D\subseteq\{-1,1\}^{n}\to\{-1,1\}$. Since the work of [beals_poly], it has been known that the acceptance probability $\mathsf{acc}(x)$ of a $T$-query bounded-error quantum algorithm on input $x\in D$ is a polynomial of degree at most $2T$. Thus, the polynomial $p(x)=1-2\cdot\mathsf{acc}(x)$ must be a good approximation of $f$. Observe that $p(x)$ remains bounded outside $D$ since $\mathsf{acc}(x)$ remains a probability defined by the algorithm, with no constraint.
@@
The following lemma connects the quantum query complexity and approximate bounded degree.
@@
Let $f:D\subseteq\{-1,1\}^{n}\to\{-1,1\}$ and $\delta>0$. If a quantum algorithm computes $f$ on $D$ with error $\delta$ using $T$ queries, then there is a polynomial $p$ of degree at most $2T$ that $2\delta$-approximates $f$ on $D$.
@@
We now consider a property of a sequence of integers as a function $F:D\subseteq[R]_{0}^{N}\to\{-1,1\}$. The symbol 00 will play a special role that will be exhibited later on. Unfortunately one cannot just take the polynomial of those integers. The standard approach (see [Aaronson02]) is to encode $s=(s_{1},\dots,s_{N})\in[R]_{0}^{N}$ into binary variables $x=(x_{i,j})_{i\in[N],j\in[R]_{0}}\in\{-1,1\}^{N(R+1)}$ encoding whether $s_{i}=j$ as follows: $x_{i,j}=-1$ if $s_{i}=j$, and $x_{i,j}=1$ otherwise. Let $H_{b}^{N(R+1)}\subseteq\{-1,1\}^{N(R+1)}$ be the set of all possible encodings of vectors $s$, that is, for every $i\in[N]$ there is exactly one $j\in[R]_{0}$ such that $x_{i,j}=-1$.
@@
But before doing this, we are going to relax the constraints on the domain $D_{b}$ in the case of symmetric functions, while we decrease its dimension. When $F$ is symmetric (i.e., $F(s)=F(s\circ\pi_{N})$ for any permutation $\pi_{N}$ of $[N]$), one can instead define a function $F_{\leq N}$ with weaker constraints by removing the variables corresponding to the symbol 00. Define $H_{\leq N}^{NR}$ as the set of length-$(NR)$ binary vectors with Hamming weight at most $N$. Given any $x\in H_{\leq N}^{NR}$, we define its frequency vector $z(x)=(z_{0},z_{1},\dots,z_{R})$ with $z_{j}=\#\{i:x_{ij}=-1\}$, for $1\leq j\leq R$, and $z_{0}=N-z_{1}-\ldots-z_{R}$. From the vector $z(x)$, one can define a valid sequence of integers $s(x)\in[R]_{0}^{N}$: it can be any sequence from $[R]_{0}^{N}$ that has frequency vector $z(x)$. Now we can define $F_{\leq N}$ on domain $D_{\leq N}$ as
@@
The following lemma connects the quantum query complexity and double-promise approximate degree.
Lemma 4.7 ([Aaronson02, ambainis_poly],[BT20, Theorem 3.9]).
Let $F:D\subseteq[R]_{0}^{N}\to\{-1,1\}$ be symmetric and $\delta>0$. Define $H_{\leq N}^{NR}\subseteq\{-1,1\}^{NR}$ and $F_{\leq N}:D_{\leq N}\subseteq H_{\leq N}^{NR}\to\{-1,1\}$ as above. If a quantum algorithm computes $F$ on $D$ with error $\delta$ using $T$ queries, then there is a polynomial $p$ of degree at most $2T$ that double-promise $2\delta$-approximates $F_{\leq N}$ on $D_{\leq N}$.